Write the equation of the line passing through each of the given points with the indicated slope. Give your results in slope-intercept form, where possible.
(0,5), m = -3/5

Answer by venugopalramana(3286)

Write the equation of the line passing through each of the given points with the indicated slope. Give your results in slope-intercept form, where possible.
(0,5), m = -3/5
STD.EQN.OF LINE THROUGH P(X1,Y1) AND SLOPE =M IS
Y-Y1=M(X-X1)
Y-5=(-3/5)(X-0)=-3X/5
Y=-3X/5 + 5